OpenStudy (anonymous):

3 to the 3rd power lus two to the 2nd power ?

OpenStudy (anonymous):

\[3^3+2^2=3\times 3\times 3+2\times 2\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[4-2\]